# Mind the Gap Why Neural Memory Fails Under Semantic Density

### v1

@@ 4.2.1 The RAG Limitation
Consider a conversational agent in which a user states “My name is Alice.” Modern vector databases can perform near-real-time upserts, so the embedding could be added quickly. However, RAG’s architectural limitations run deeper than indexing latency. First, RAG is append-first: adding the new fact does not remove or supersede any prior statement about the user’s name, so both old and new values may be retrieved. Second, RAG lacks key-based identity: there is no canonical address for “the user’s name” that can be surgically updated; facts are identified only by embedding similarity, which is fuzzy. Third, RAG provides no schema guarantees: the fact might be stored as “user’s name is Alice” or “Alice is the user’s name” depending on phrasing, breaking downstream systems that expect consistent structure. Fourth, RAG has no native version semantics: if the user later says “Actually, call me Alicia,” the system has no mechanism to mark the prior value as superseded—both will be retrieved with similar scores.
These limitations are not addressed by standard similarity-only RAG; resolving them requires adding explicit identity, schema, and version semantics—effectively introducing KO-like structure. Knowledge Objects provide these capabilities natively: hash-based addressing provides key-based identity for surgical updates, controlled vocabularies enforce schema consistency, version chains provide explicit supersession semantics, and the combination enables true online episodic memory where facts are available immediately with deterministic overwrite behavior.
@@ 4.2.2 Knowledge Object Data Structure
4.2.3 Storage: Hash-Addressed with $O(1)$ Writes
Knowledge Objects are stored in a hash map keyed by $\text{hash}(\text{subject}\|\text{predicate})$, which provides constant-time insertion and update, constant-time exact lookup when the subject and predicate are known, and zero interference between facts since distinct keys map to distinct addresses and updating one fact cannot corrupt another. This last property matters most: whereas neural memory systems store all facts in shared continuous parameters where updates necessarily interact, hash-based storage guarantees complete isolation between memory entries.
@@ 4.2.4 Retrieval: Semantic Search over Embeddings
For queries that do not specify an exact subject-predicate pair, we perform $k$-nearest neighbor search over the embedding space. The process begins by encoding the query as $\mathbf{q}=\text{Encoder}(\text{query})$, then finding the top-$k$ Knowledge Objects by cosine similarity via $\text{argmax}_{k}\cos(\mathbf{q},\mathbf{e}_{i})$, and finally returning the matched KOs to augment the LLM’s response. This hybrid approach—exact lookup when possible, semantic search otherwise—combines the precision of structured databases with the flexibility of neural retrieval.
@@ 4.2.5 The KO Compiler: Extracting Structured Facts
Raw text must be compiled into structured Knowledge Objects, a process we accomplish through an LLM-based compiler. The compiler receives natural language input and extracts factual statements as subject-predicate-object triples, using snake_case for predicates to ensure consistency and including confidence scores for each extraction. For instance, given the input “Dr. Sarah Chen is a 42-year-old cardiologist at MGH. Her phone is 617-555-0123,” the compiler produces four triples: (Dr_Sarah_Chen, age, 42), (Dr_Sarah_Chen, occupation, cardiologist), (Dr_Sarah_Chen, employer, MGH), and (Dr_Sarah_Chen, phone_number, 617-555-0123).
@@ 4.2.6 Surgical Updates: $O(1)$ Correction
When a fact changes, KO performs a surgical update by computing the identifier $\text{id}=\text{hash}(\text{subject}\|\text{predicate})$, optionally archiving the old value with a version pointer for history preservation, and writing the new value at the same address. This operation completes in $O(1)$ time regardless of corpus size. By contrast, RAG systems must append the new fact to a pending queue, re-index the entire corpus or affected partition, and handle the conflict period during which old and new facts coexist until re-indexing completes—a process that scales with corpus size and introduces temporal inconsistency.
@@ 4.2.7 Version Chains: Temporal Queries
Knowledge Objects optionally preserve version history, maintaining a linked list of prior values with timestamps. For example, a user’s favorite food might have evolved through versions: spaghetti (January 2024, archived), sushi (June 2024, archived), and pizza (January 2025, current). This structure enables temporal queries: asking “What is my favorite food?” returns the current value (pizza), asking “What was my favorite food before?” returns the previous version (sushi), and asking “What changed since January?” triggers traversal of the version chain. Neural memory fundamentally cannot support such queries because weight updates destroy prior states irrecoverably—there is no “undo” operation for gradient descent.
@@ 4.2.9 Illustrative Example: A Medical Research Assistant
To make the architecture concrete, consider a medical research assistant that helps a physician track patient cases over months of interaction. In a first session during January, the physician states: “Patient M.R., 67-year-old male, presented with persistent cough and 15-pound weight loss. CT scan showed 3cm mass in right upper lobe. Biopsy confirmed non-small cell lung cancer, stage IIB.” The KO compiler extracts structured facts—(patient_MR, age, 67), (patient_MR, diagnosis, NSCLC_IIB), (patient_MR, tumor_size, 3cm), and (patient_MR, tumor_location, right_upper_lobe)—which are written immediately to the KO store and become available for retrieval in subsequent queries without waiting for batch reindexing.
@@ 4.3.3 The Learned Router
We experimented with two router variants: (1) logistic regression on raw query embeddings (384-dimensional vectors from the sentence transformer), and (2) logistic regression on handcrafted surface features extracted from the query text. Both approaches use a training set of 30 examples comprising 15 factual queries (such as “What is X?” and “Who created Y?”) and 15 fuzzy queries (such as “Write a poem” and “Explain the concept”). The surface-feature variant, which we report in our experiments, extracts features such as: presence of interrogative words (what, who, when, where), question mark termination, proper noun indicators, word count, and creative/explanatory verbs (write, help, suggest, explain). This variant achieved slightly better accuracy and interpretability, though both approaches exceeded 95% accuracy. We use logistic regression as the simplest router; in Experiment 5 (Table 6) we also evaluate other lightweight classifiers (Random Forest, SVM, Gradient Boosting) on the same feature set to show that the routing task is robustly learnable. What makes this work is that the router examines linguistic structure rather than embedding similarity scores, making it immune to vector space crowding: a query like “What is the capital of France?” is classified as factual regardless of how crowded the embedding space becomes.
Figure 3: Hybrid Architecture. Queries are embedded and classified by a Learned Router. Factual queries retrieve from the KO hash map; fuzzy queries pass directly to the LLM. This division of labor—discrete memory for facts, neural weights for generation—is the core architectural principle.
@@ 5.1 Experiment 1: The Stability Gap
We evaluated both memory systems on synthetic entity-attribute-value facts following the template “What is the [attribute] of [entity]?” with answers drawn from a controlled vocabulary. The synthetic corpus was designed to exhibit high semantic density ($\rho\approx 0.70$), representing a challenging but realistic scenario where facts share similar linguistic structure—as would occur in a database of customer records, patient information, or product specifications where queries follow consistent patterns. We varied the corpus size $N$ from 1 to 75 facts in increments chosen to reveal the collapse trajectory. For each value of $N$, we stored all facts in both memory systems, then queried each fact and measured retrieval accuracy as the percentage of queries returning the correct answer. All experiments were run for 5 independent trials with different random seeds, and we report mean accuracy. The neural baseline (Linearized Attention Memory) was configured with embedding dimension $d=384$ using the all-MiniLM-L6-v2 sentence transformer, decay factor $\lambda=0.99$, and learning rate $\eta=0.1$. The discrete baseline (Knowledge Objects) used the same embedding model for semantic search, with facts stored in a hash map keyed by subject-predicate pairs.
Table 2: Stability Gap Results: Neural collapse under high semantic density ($\rho\approx 0.70$). Gap measured in percentage points ($p$).
@@
Neural memory achieves 100% accuracy at $N=1$, confirming correct implementation of the storage and retrieval mechanisms. As $N$ increases, accuracy collapses rapidly: 40% at $N=5$, 5% at $N=20$, and 1.3% at N=75N=75—barely above the random baseline of $1/75\approx 1.3\%$. Discrete memory maintains 100% accuracy throughout the entire range, yielding a stability gap of 98.7 percentage points at $N=75$. The high semantic density ($\rho\approx 0.70$) accelerates collapse compared to corpora with more diverse fact structures; this confirms that semantic density $\rho$ is a key predictor of neural memory failure, with higher density producing faster collapse.
Figure 4: The Stability Gap. Neural memory (red) collapses as NN increases, while discrete memory (green) maintains 100% accuracy throughout. Embedding dimension $d=384$ is held constant; only the number of stored facts varies. The shaded region represents the accuracy gap caused by write-time interference in continuous vector storage. Even as we store more facts, no improvement in attention or retrieval mechanism can prevent this collapse—it is a geometric consequence of semantic overlap in shared parameters.
@@
To further validate the relationship between semantic density and collapse, we generated fact corpora with varying $\rho$ values by manipulating the diversity of entity names, attribute types, and question structures. Figure 5 plots $N_{50}$ (the number of facts at which neural accuracy drops to 50%) against semantic density. The relationship follows an exponential decay: at $\rho\approx 0.34$, collapse occurs around $N=20$; at $\rho>0.6$, collapse is nearly immediate with $N_{50}\leq 5$. We could not generate realistic fact corpora with $\rho<0.3$—even maximally diverse templates (different domains, question structures, entity types) produce baseline semantic similarity because natural language has inherent structure that embeddings capture.
@@ 5.3 Experiment 3: Deterministic Overwrite Semantics (KO vs RAG)
While modern vector databases support near-real-time upserts, RAG’s architectural limitations extend beyond indexing latency. RAG is fundamentally append-first and retrieval-first: new facts are added to the index, but there is no mechanism to identify and remove the prior value for the same logical fact. If a user says “My phone is 555-1111” and later “My phone changed to 555-2222,” both facts exist in the index with similar embeddings, and retrieval may return either or both depending on query phrasing and similarity scores. This is not a bug—it reflects RAG’s design for static corpora where facts accumulate rather than mutate.
@@
We simulated a conversational session with 6 fact-learning exchanges, each following the same pattern: the user provides a fact in natural language (such as “My name is Alice” or “My birthday is March 15th”), the system extracts and stores the fact, and then we immediately query for that fact using a differently-phrased question (such as “What is my name?” or “When is my birthday?”). The key measurement is whether the system can retrieve the just-learned fact with deterministic accuracy. Our RAG baseline models a common production deployment pattern where new documents enter a pending queue for batch indexing—while some vector databases support near-real-time upserts, many production RAG systems use asynchronous ingestion pipelines for throughput and cost reasons. We also measured accuracy after a simulated batch reindex to establish an upper bound on eventual performance.
@@
Knowledge Objects retrieve newly-learned facts immediately with 100% accuracy because hash-based storage provides key-based identity: the fact “user’s name” has a deterministic address that is overwritten when updated. The batch-ingestion RAG baseline returns 0% before reindexing because new embeddings have not yet been added to the searchable index. Even after reindexing, RAG achieves only 83.3% accuracy due to embedding mismatch between the phrasing of statements (“My birthday is March 15th”) and queries (“When is my birthday?”), a known limitation of embedding-based retrieval when statement and query forms differ substantially. We note that RAG systems with real-time upsert capabilities would eliminate the “immediate retrieval” gap, but would still lack the key-based identity and version semantics needed for deterministic overwrite—both old and new values would coexist in the index with similar embeddings.
@@ 5.4 Experiment 4: Surgical Correction
Knowledge Objects achieve perfect surgical updates: 100% of updated facts return the new value, and 100% of other facts remain unchanged, with each update completing in $O(1)$ time regardless of corpus size. Neural memory exhibits catastrophic forgetting: 0% of updates succeed (the system returns neither old nor new values reliably), and 98.9% of other facts are corrupted by the update operation. This occurs because gradient-based updates to shared weight matrices necessarily affect all stored associations, not just the target fact. Discrete addressing provides complete isolation between facts—updating one cannot affect another—while neural memory fundamentally cannot support surgical correction without catastrophic interference.

### v2

@@ 4.2.1 The RAG Limitation
Consider a conversational agent in which a user states “My name is Alice.” Modern vector databases can perform near-real-time upserts, so the embedding could be added quickly. However, RAG’s architectural limitations run deeper than indexing latency. First, RAG is append-first: adding the new fact does not remove or supersede any prior statement about the user’s name, so both old and new values may be retrieved. Second, RAG lacks key-based identity: there is no canonical address for “the user’s name” that can be surgically updated; facts are identified only by embedding similarity, which is fuzzy. Third, RAG provides no schema guarantees: the fact might be stored as “user’s name is Alice” or “Alice is the user’s name” depending on phrasing, breaking downstream systems that expect consistent structure. Fourth, RAG has no native version semantics: if the user later says “Actually, call me Alicia,” the system has no mechanism to mark the prior value as superseded, both will be retrieved with similar scores.
These limitations are not addressed by standard similarity-only RAG; resolving them requires adding explicit identity, schema, and version semantics, effectively introducing KO-like structure. Knowledge Objects provide these capabilities natively: hash-based addressing provides key-based identity for surgical updates, controlled vocabularies enforce schema consistency, version chains provide explicit supersession semantics, and the combination enables true online episodic memory where facts are available immediately with deterministic overwrite behavior.
@@ 4.2.2 Knowledge Object Data Structure
The Knowledge Object architecture supports richer content structures including multi-sentence claims, hypotheses, and reasoning traces; we use subject-predicate-object facts here to isolate the retrieval mechanism from content complexity.
4.2.3 Storage: Pattern Separation via Hash Addressing
Knowledge Objects are stored in a hash map keyed by $\text{hash}(\text{subject}\|\text{predicate})$. This hash function implements pattern separation: it takes semantically overlapping inputs (embedding vectors with $\cos\theta>0$) and maps them to discrete addresses (which are orthogonal by construction, with overlap exactly 0). The resulting storage provides constant-time insertion and update, constant-time exact lookup when the subject and predicate are known, and zero interference between facts since distinct keys map to distinct addresses and updating one fact cannot corrupt another. This last property matters most: whereas neural memory systems store all facts in shared continuous parameters where updates necessarily interact, hash-based storage guarantees complete isolation between memory entries, the same guarantee the Dentate Gyrus provides for biological episodic memory.
@@ 4.2.4 Retrieval: Semantic Search over Embeddings
For queries that do not specify an exact subject-predicate pair, we perform $k$-nearest neighbor search over the embedding space. The process begins by encoding the query as $\mathbf{q}=\text{Encoder}(\text{query})$, then finding the top-$k$ Knowledge Objects by cosine similarity via $\text{argmax}_{k}\cos(\mathbf{q},\mathbf{e}_{i})$, and finally returning the matched KOs to augment the LLM’s response. This hybrid approach, exact lookup when possible, semantic search otherwise, combines the precision of structured databases with the flexibility of neural retrieval.
4.2.5 Two Retrieval Modes: A Critical Distinction
Knowledge Objects support two fundamentally different retrieval modes with distinct accuracy and latency characteristics, and conflating them leads to confusion about system capabilities.
Hash-based retrieval applies when the subject and predicate are known (e.g., a structured query for (France, capital, ?)). For predicates that are functional, where each subject has exactly one value, hash lookup achieves 100% accuracy with $O(1)$ latency. Multi-valued predicates (a country has multiple neighbors, a person has multiple employers over time) require standard extensions such as list storage or composite keys.
Embedding-based retrieval applies when queries are natural language (e.g., “What is the capital of France?”). This mode relies on the embedding model’s ability to map the query near the correct fact in vector space. Accuracy depends on embedding quality and semantic density of the corpus, we observe approximately 46% top-1 accuracy on Wikipedia facts, with 78% top-5 recall. Latency scales as $O(N)$ for brute-force search or $O(\log N)$ with approximate nearest neighbor indices such as HNSW.
The experiments in Section 5 primarily evaluate embedding-based retrieval because it represents the harder case: users issue natural language queries, and the system must locate relevant facts without knowing the exact keys. Production deployments can improve effective accuracy through query parsing (converting natural language to structured queries when possible), multi-value hash storage, reranking (using a cross-encoder to rescore top-$k$ candidates), or hybrid routing (attempting hash lookup first, falling back to embedding search).
@@ 4.2.6 The KO Compiler: Extracting Structured Facts
Raw text must be compiled into structured Knowledge Objects, a process we accomplish through an LLM-based compiler. The compiler receives natural language input and extracts factual statements as subject-predicate-object facts, using snake_case for predicates to ensure consistency and including confidence scores for each extraction. For instance, given the input “Dr. Sarah Chen is a 42-year-old cardiologist at MGH. Her phone is 617-555-0123,” the compiler produces four facts: (Dr_Sarah_Chen, age, 42), (Dr_Sarah_Chen, occupation, cardiologist), (Dr_Sarah_Chen, employer, MGH), and (Dr_Sarah_Chen, phone_number, 617-555-0123).
This extraction step is where schema drift can occur: without controlled vocabularies constraining predicate choices, an LLM might extract the same fact as phone_number, phone, or contact_phone depending on phrasing. Experiment 7 quantifies this drift across frontier models. Our retrieval experiments (Experiments 1–6, 8) use pre-structured data (Wikipedia facts, templated scientific measurements) to isolate the retrieval mechanism from extraction variability, this represents the best-case scenario where extraction is perfect, establishing an upper bound on system performance.
@@ 4.2.7 Surgical Updates: $O(1)$ Correction
When a fact changes, KO performs a surgical update by computing the identifier $\text{id}=\text{hash}(\text{subject}\|\text{predicate})$, optionally archiving the old value with a version pointer for history preservation, and writing the new value at the same address. This operation completes in $O(1)$ time regardless of corpus size. By contrast, RAG systems must append the new fact to a pending queue, re-index the entire corpus or affected partition, and handle the conflict period during which old and new facts coexist until re-indexing completes, a process that scales with corpus size and introduces temporal inconsistency.
@@ 4.2.8 Version Chains: Temporal Queries
Knowledge Objects optionally preserve version history, maintaining a linked list of prior values with timestamps. For example, a user’s favorite food might have evolved through versions: spaghetti (January 2024, archived), sushi (June 2024, archived), and pizza (January 2025, current). This structure enables temporal queries: asking “What is my favorite food?” returns the current value (pizza), asking “What was my favorite food before?” returns the previous version (sushi), and asking “What changed since January?” triggers traversal of the version chain. Neural memory cannot readily support such queries because weight updates destroy prior states irrecoverably, there is no “undo” operation for gradient descent.
@@ 4.2.10 Illustrative Example: A Medical Research Assistant
To make the architecture concrete, consider a medical research assistant that helps a physician track patient cases over months of interaction. In a first session during January, the physician states: “Patient M.R., 67-year-old male, presented with persistent cough and 15-pound weight loss. CT scan showed 3cm mass in right upper lobe. Biopsy confirmed non-small cell lung cancer, stage IIB.” The KO compiler extracts structured facts, (patient_MR, age, 67), (patient_MR, diagnosis, NSCLC_IIB), (patient_MR, tumor_size, 3cm), and (patient_MR, tumor_location, right_upper_lobe), which are written immediately to the KO store and become available for retrieval in subsequent queries without waiting for batch reindexing.
@@ 4.3.3 The Learned Router
We experimented with two router variants: (1) logistic regression on raw query embeddings (384-dimensional vectors from the sentence transformer), and (2) logistic regression on handcrafted surface features extracted from the query text. Both approaches use a training set of 30 examples comprising 15 factual queries (such as “What is X?” and “Who created Y?”) and 15 fuzzy queries (such as “Write a poem” and “Explain the concept”). We include the router to demonstrate feasibility of intent separation; we do not claim this specific classifier is optimal, and production deployments would likely benefit from larger training sets and more sophisticated classifiers. The surface-feature variant, which we report in our experiments, extracts features such as: presence of interrogative words (what, who, when, where), question mark termination, proper noun indicators, word count, and creative/explanatory verbs (write, help, suggest, explain). This variant achieved slightly better accuracy and interpretability, though both approaches exceeded 95% accuracy. We use logistic regression as the simplest router; in Experiment 5 (Table 6) we also evaluate other lightweight classifiers (Random Forest, SVM, Gradient Boosting) on the same feature set to show that the routing task is robustly learnable. What makes this work is that the router examines linguistic structure rather than embedding similarity scores, making it immune to vector space crowding: a query like “What is the capital of France?” is classified as factual regardless of how crowded the embedding space becomes.
Figure 3: Hybrid Architecture. Queries are embedded and classified by a Learned Router. Factual queries retrieve from the KO hash map; fuzzy queries pass directly to the LLM. This division of labor, discrete memory for facts, neural weights for generation, is the core architectural principle.
@@ 5.1 Experiment 1: Semantic Interference
We evaluated both memory systems on synthetic entity-attribute-value facts following the template “What is the [attribute] of [entity]?” with answers drawn from a controlled vocabulary. The synthetic corpus was designed to exhibit high semantic density ($\rho\approx 0.70$), representing a challenging but realistic scenario where facts share similar linguistic structure, as would occur in a database of customer records, patient information, or product specifications where queries follow consistent patterns. We varied the corpus size $N$ from 1 to 75 facts in increments chosen to reveal the collapse trajectory. For each value of $N$, we stored all facts in both memory systems, then queried each fact and measured retrieval accuracy as the percentage of queries returning the correct answer. All experiments were run for 5 independent trials with different random seeds, and we report mean accuracy. The neural baseline (Linearized Attention Memory) was configured with embedding dimension $d=384$ using the all-MiniLM-L6-v2 sentence transformer, decay factor $\lambda=0.99$, and learning rate $\eta=0.1$. The discrete baseline (Knowledge Objects) used the same embedding model for semantic search, with facts stored in a hash map keyed by subject-predicate pairs.
Table 2: Semantic Interference Results: Neural collapse under high semantic density ($\rho\approx 0.70$). Gap measured in percentage points ($p$).
@@
Neural memory achieves 100% accuracy at $N=1$, confirming correct implementation of the storage and retrieval mechanisms. As $N$ increases, accuracy collapses rapidly: 40% at $N=5$, 5% at $N=20$, and 1.3% at $N=75$, barely above the random baseline of $1/75\approx 1.3\%$. Discrete memory maintains 100% accuracy throughout the entire range, yielding an accuracy gap of 98.7 percentage points at $N=75$. The high semantic density ($\rho\approx 0.70$) accelerates collapse compared to corpora with more diverse fact structures; this confirms that semantic density $\rho$ is a key predictor of neural memory failure, with higher density producing faster collapse.
Figure 4: Semantic Interference. Neural memory (red) collapses as NN increases, while discrete memory (green) maintains 100% accuracy throughout. Embedding dimension $d=384$ is held constant; only the number of stored facts varies. The shaded region represents the accuracy gap caused by write-time interference in continuous vector storage. Even as we store more facts, no improvement in attention or retrieval mechanism can prevent this collapse: it is a geometric consequence of semantic overlap in shared parameters.
@@
To further validate the relationship between semantic density and collapse, we generated fact corpora with varying $\rho$ values by manipulating the diversity of entity names, attribute types, and question structures. Figure 5 plots $N_{50}$ (the number of facts at which neural accuracy drops to 50%) against semantic density. The relationship follows an exponential decay: at $\rho\approx 0.34$, collapse occurs around $N=20$; at $\rho>0.6$, collapse is nearly immediate with $N_{50}\leq 5$. We could not generate realistic fact corpora with $\rho<0.3$, even maximally diverse templates (different domains, question structures, entity types) produce baseline semantic similarity because natural language has inherent structure that embeddings capture.
@@ 5.3 Experiment 3: Deterministic Overwrite Semantics (KO vs RAG)
While modern vector databases support near-real-time upserts, RAG’s architectural limitations extend beyond indexing latency. RAG is fundamentally append-first and retrieval-first: new facts are added to the index, but there is no mechanism to identify and remove the prior value for the same logical fact. If a user says “My phone is 555-1111” and later “My phone changed to 555-2222,” both facts exist in the index with similar embeddings, and retrieval may return either or both depending on query phrasing and similarity scores. This is not a bug, it reflects RAG’s design for static corpora where facts accumulate rather than mutate.
@@
We simulated a conversational session with 6 fact-learning exchanges, each following the same pattern: the user provides a fact in natural language (such as “My name is Alice” or “My birthday is March 15th”), the system extracts and stores the fact, and then we immediately query for that fact using a differently-phrased question (such as “What is my name?” or “When is my birthday?”). The key measurement is whether the system can retrieve the just-learned fact with deterministic accuracy. Our RAG baseline models a common production deployment pattern where new documents enter a pending queue for batch indexing, while some vector databases support near-real-time upserts, many production RAG systems use asynchronous ingestion pipelines for throughput and cost reasons. We also measured accuracy after a simulated batch reindex to establish an upper bound on eventual performance.
@@
Knowledge Objects retrieve newly-learned facts immediately with 100% accuracy because hash-based storage provides key-based identity: the fact “user’s name” has a deterministic address that is overwritten when updated. The batch-ingestion RAG baseline returns 0% before reindexing because new embeddings have not yet been added to the searchable index. Even after reindexing, RAG achieves only 83.3% accuracy due to embedding mismatch between the phrasing of statements (“My birthday is March 15th”) and queries (“When is my birthday?”), a known limitation of embedding-based retrieval when statement and query forms differ substantially. We note that RAG systems with real-time upsert capabilities would eliminate the “immediate retrieval” gap, but would still lack the key-based identity and version semantics needed for deterministic overwrite, both old and new values would coexist in the index with similar embeddings.
@@ 5.4 Experiment 4: Surgical Correction
Knowledge Objects achieve perfect surgical updates: 100% of updated facts return the new value, and 100% of other facts remain unchanged, with each update completing in $O(1)$ time regardless of corpus size. Neural memory exhibits catastrophic forgetting: 0% of updates succeed (the system returns neither old nor new values reliably), and 98.9% of other facts are corrupted by the update operation. This occurs because gradient-based updates to shared weight matrices necessarily affect all stored associations, not just the target fact. Discrete addressing provides complete isolation between facts, updating one cannot affect another, while neural memory fundamentally cannot support surgical correction without catastrophic interference.